27. A classroom that normally contains 40 people is to be air-conditioned with window air-conditioning units of 5 KW cooling capacity. A person at rest may be assumed to dissipate heat at a rate of about 360 KJ/hr. There are 10 light bulbs in the room, each with a rating of 100 watts. The rate of heat transfer to the classroom through the walls and the windows is estimated to be 15,000 KJ/hr. If the room air is to be maintained at a constant temperature of 21oC, determine the number of window air-conditioning units required.


1
Expert's answer
2022-03-13T15:15:18-0400

We assume that there is no heat dissipating instrument in the room

Total cooling load of the room is defined from the given below equation

Qcooling = Qlight + Qpeople + Qheatgain

where

Qlight = 10*100 W = 1 KW

Qpeople = 40*360 KJ/h = 4 KW

Qheatgain = 15000KJ/h = 4.17 KW

Qcooling = 1 + 4 + 4.17 = 9.17 KW

the number of air conditioner unit is = 9.17/5 = 1.83, which is approximately 2 units